OpenStudy (anonymous):

Select the equation that most accurately depicts the word problem. A class of 19 pupils has five more girls than boys. Let n = the number of boys. n + (n + 19) = 5 n + (n + 5) = 19 n - (n + 5) = 19 n + (n - 19) = 5

OpenStudy (gorv):

so boys =n girls=no.of boys+5=n+5 total=19=n+n+5